OpenStudy (anonymous):

1. add or subtract 3/x-3-5/x-2 2.add or subtract -5/x-3/-4/x+2

OpenStudy (pfenn1):

1.Is this the correct expression?\[-\frac{4x}{x+7}-\frac{8}{x+7} \]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

\[-\frac{4x}{x+7}-\frac{8}{x+7}=\frac{-4x-8}{x+7}\]

OpenStudy (anonymous):

thanks

OpenStudy (pfenn1):

Have you tried #2?

OpenStudy (anonymous):

no

OpenStudy (pfenn1):

2. \[\frac{3}{x-3}-\frac{5}{x-2}\]This is similar, but in order to continue, the denominators must be the same. What is the LCD?

OpenStudy (anonymous):

i dont know

OpenStudy (pfenn1):

Well, here is a site that helps to understand the concept of LCD. In this case I think I would use (x-3)(x-2) as the LCD.\[\left( \frac{x-2}{x-2} \right)\frac{3}{x-3}-\left( \frac{x-3}{x-3} \right)\frac{5}{x-2}=\frac{3(x-2)-5(x-3)}{(x-2)(x-3)}\]Can you take it from here?

OpenStudy (pfenn1):

*here is the website: http://www.mathsisfun.com/least-common-denominator.html